OpenStudy (heart_offire):

2x + 4y = 12 3x - 5y = 10 Question: A. Write the equations in slope-intercept form. (Show your work.) B. Graph the pair of linear equations. C. Use the graph to estimate the solution to the system of equations.

OpenStudy (mathmale):

2x + 4y = 12 we want to solve this for y, and want the final result to look like y=mx+b. How would you go about isolating 4y on the left side? Please share your work.

OpenStudy (heart_offire):

well i looked up how to convert standard for to slope intercept.  this is for 2x +4y=12